### v2

@@ 5 Conclusions
This study provides a comprehensive empirical assessment of the forecast stability of global models, focusing on the effects of retraining across both point and probabilistic forecasting, in the context of retail demand. Our findings demonstrate that global forecasting models are able to produce very stable predictions, and that they can achieve even more stability under reduced retraining strategies, challenging the prevailing wisdom that frequent model updates are necessary to maintain forecast stability.
@@
In terms of point forecasting, the results consistently show that retraining less frequently does not harm and often improves forecast stability. This trend is evident in both the M5 and the VN1 datasets, where many models exhibit monotonic or flat stability profiles as the retraining interval increases. Probabilistic forecast stability, as measured by the newly defined Scaled Multi-Quantile Change (SMQC), shows a very similar behavior, indicating that the stability of probabilistic forecasts does not deteriorate as the models are retrained less frequently. These results are statistically validated via Friedman-Nemenyi tests and indicate that continuous retraining is often not the optimal strategy.
Ensembling can be considered a good strategy to enhance forecast stability. Notably, ensemble models not only improve stability metrics but also smooth the variation in stability across retraining scenarios, further reinforcing their role as a stabilizing force in global forecasting systems. However, from our analysis, it is not clear whether the size and the diversity of the ensemble may have an impact on stability, and more research is needed in this direction.
The joint analysis of retraining and ensembling reveals that the two approaches act as complementary levers for stabilizing forecasts. Less frequent retraining reduces abrupt changes in model behavior across forecasting origins, while ensembling mitigates the instability arising from any single model’s sensitivity to input changes. When combined, these strategies offer a robust forecasting setup that minimizes the instability of both point estimates and prediction intervals without sacrificing predictive accuracy. Figures comparing the accuracy–stability trade-off clearly show that ensembles lie closer to the optimality, offering an optimal balance between these competing objectives.
@@
From a practical standpoint, these findings have several implications. First, businesses that rely on large-scale forecasting systems, such as those in retail, inventory management, and supply chain planning can adopt lower retraining strategies not only to conserve computational resources, but also to increase operational stability, without compromising on forecast accuracy. Second, incorporating ensembles into global forecasting frameworks provides an effective hedge against model instability, thereby increasing the reliability of forecasts used in decision-making processes. Finally, these results support a shift in emphasis from pure accuracy to stability-aware model selection and evaluation, in those contexts that are more sensitive to forecast reviews.
@@
Nevertheless, some limitations of the present study should be acknowledged. While the analysis covers a broad range of global models and two state-of-the-art retail forecasting datasets, the generalizability of the findings to other domains or data types (e.g., macroeconomic indicators, financial series, or healthcare time series) remains an open question. Moreover, this work deliberately focuses on the stability of global models, leaving the analysis of stability in local models as an open avenue. A direct comparison between global and local approaches could therefore represent a valuable direction for future research. Additionally, the study focuses solely on full retraining schemes and does not explore incremental or online learning strategies, which may offer a different trade-off between accuracy and stability. Future research could extend this work by incorporating such adaptive methods, assessing alternative ensembling techniques (e.g., weighted or stacking ensembles), and introducing business-oriented stability metrics that link forecast instability directly to operational cost or risk. Furthermore, this study employed a single framework for generating probabilistic forecasts, namely Conformal Inference. While conformalized residuals provide a valid means for quantifying uncertainty, the resulting quantiles may primarily capture the stability of the conformal adjustment rather than that of the underlying predictive distributions. Future research could address this limitation by studying probabilistic stability in native probabilistic models, thereby enhancing the robustness and generalizability of the findings. Finally, further theoretical development of stability measures, particularly for probabilistic forecasts, may also contribute to a deeper understanding of how models behave across time and under varying data regimes.
In conclusion, this work provides strong empirical evidence that frequent retraining is not necessary for maintaining stability in global forecasting models, and that ensembling is an effective strategy for achieving both stable and accurate forecasts. These findings contribute to a growing body of research promoting sustainable and robust forecasting practices in modern data-intensive environments.
@@
In this section, we provide tables and figures related to the empirical results of the M5 and VN1 datasets.
Table 3 shows the composition of ensembles across the different datasets.
The Tables 4, 5, 7, and 6 show the forecast instability of the different models and ensembles along the examined retrain scenarios for the M5 daily dataset.
The Tables 8, 9, 11, and 10 show the forecast instability of the different models and ensembles along the examined retrain scenarios for the VN1 weekly dataset.
Figures 9, and 10 show the results of the Friedman-Nemenyi test in the context of probabilistic instability for the M5 and VN1 datasets.
Figures 11, 12, 13, and 14 show the results of the Friedman-Nemenyi test in the context of both point and probabilistic instability for the different ensemble models across retrain scenarios and for each dataset.
Table 3: Composition of ensembles for the M5 and VN1 datasets.
Dataset Ensemble Models Included
M5 ENS2A XGBoost, LGBM
ENS3A XGBoost, LGBM, LR
ENS4A XGBoost, LGBM, LR, NBEATSx
ENS5A XGBoost, LGBM, LR, NBEATSx, MLP
VN1 ENS2A MLP, NBEATSx
ENS3A MLP, NBEATSx, NHITS
ENS4A MLP, NBEATSx, NHITS, RF
ENS5A MLP, NBEATSx, NHITS, RF, XGBoost
@@
LR 0.188 0.188 0.188 0.189 0.189 0.190 0.190 0.192 0.193 0.196
XGBoost 0.075 0.074 0.073 0.073 0.072 0.072 0.072 0.072 0.073 0.073
@@
CatBoost 0.118 0.043 0.059 0.037 0.029 0.020 0.018 0.018 0.018 0.019
MLP 0.500 0.511 0.506 0.520 0.513 0.524 0.510 0.525 0.475 0.553
TCN 0.499 0.209 0.174 0.086 0.055 0.043 0.031 0.017 0.017 0.003
NBEATSx 0.547 0.541 0.527 0.561 0.528 0.536 0.523 0.536 0.515 0.412
NHITS 0.566 0.548 0.558 0.534 0.565 0.537 0.533 0.520 0.512 0.564
@@
Ens2A 0.059 0.058 0.058 0.058 0.058 0.058 0.058 0.058 0.058 0.058
Ens3A 0.089 0.089 0.089 0.089 0.089 0.089 0.090 0.090 0.091 0.092
Ens4A 0.095 0.094 0.094 0.097 0.095 0.095 0.095 0.095 0.096 0.097
Ens5A 0.099 0.098 0.098 0.101 0.099 0.099 0.099 0.099 0.100 0.101
Table 5: M5 sMAPC values for each ensemble and retrain scenario combination.
@@
Method 7 14 21 30 60 90 120 150 180 364
Ens2A 0.108 0.108 0.108 0.108 0.108 0.108 0.108 0.108 0.108 0.107
Ens3A 0.107 0.107 0.107 0.108 0.107 0.107 0.107 0.108 0.107 0.107
Ens4A 0.099 0.099 0.099 0.101 0.100 0.100 0.099 0.100 0.099 0.099
Ens5A 0.095 0.095 0.095 0.096 0.096 0.095 0.095 0.096 0.095 0.095
Table 7: M5 SMQC values for each ensemble and retrain scenario combination.
@@
LR 0.085 0.084 0.085 0.082 0.083 0.086 0.086 0.085 0.084 0.082
RF 0.092 0.090 0.089 0.089 0.088 0.088 0.089 0.089 0.089 0.090
XGBoost 0.146 0.129 0.126 0.120 0.118 0.115 0.114 0.119 0.117 0.113
@@
CatBoost 0.320 0.198 0.186 0.152 0.134 0.139 0.115 0.117 0.100 0.095
MLP 0.458 0.469 0.468 0.475 0.443 0.483 0.438 0.490 0.469 0.452
LSTM 0.558 0.301 0.203 0.126 0.101 0.044 0.055 0.048 0.032 0.013
TCN 0.600 0.260 0.226 0.177 0.120 0.091 0.052 0.056 0.020 0.013
NBEATSx 0.596 0.594 0.596 0.577 0.577 0.581 0.573 0.580 0.571 0.550
NHITS 0.605 0.589 0.583 0.572 0.533 0.541 0.552 0.528 0.573 0.580
@@
Method 7 14 21 30 60 90 120 150 180 364
Ens2A 0.503 0.515 0.505 0.495 0.499 0.504 0.487 0.485 0.495 0.509
Ens3A 0.514 0.526 0.526 0.509 0.509 0.514 0.510 0.512 0.511 0.512
Ens4A 0.167 0.170 0.162 0.153 0.145 0.137 0.148 0.134 0.135 0.145
Ens5A 0.164 0.162 0.154 0.145 0.138 0.131 0.138 0.129 0.129 0.135
Table 9: VN1 sMAPC values for each ensemble and retrain scenario combination.
@@
LR 1.560 1.552 1.546 1.543 1.562 1.525 1.547 1.643 1.682 1.900
@@
Method 7 14 21 30 60 90 120 150 180 364
Ens2A 0.441 0.476 0.452 0.391 0.354 0.332 0.364 0.300 0.299 0.312
Ens3A 0.539 0.562 0.507 0.430 0.389 0.324 0.398 0.312 0.309 0.322
Ens4A 0.545 0.557 0.522 0.465 0.433 0.387 0.443 0.376 0.376 0.389
Ens5A 0.540 0.545 0.519 0.472 0.448 0.410 0.453 0.407 0.405 0.416
Table 11: VN1 SMQC values for each ensemble and retrain scenario combination.
@@
Figure 11: M5 Friedman-Nemenyi test results on ensembles based on sMAPC.
Figure 12: M5 Friedman-Nemenyi test results on ensembles based on SMQC.
Figure 13: VN1 Friedman-Nemenyi test results on ensembles based on sMAPC.
Figure 14: VN1 Friedman-Nemenyi test results on ensembles based on SMQC.

### v3

@@ 5 Conclusions
This study investigates the role of retraining frequency in global forecasting models, with a particular focus on its implications for both forecast accuracy and forecast stability in point and probabilistic settings. Using three large-scale benchmark datasets (M4, M5, and VN1), we provided an extensive empirical analysis across multiple machine learning and deep learning architectures, complemented by a comparison with classical local forecasting methods. These results contribute to a more nuanced understanding of retraining as a design choice in large-scale forecasting systems, rather than as a default operational requirement.
@@
Across all datasets and modeling families, we find that reducing the retraining frequency does not harm and often improves forecast stability. This result holds for both point and probabilistic forecasts, as measured by the proposed Scaled Multi-Quantile Change (SMQC) metric. In particular, stability profiles are typically flat or monotonically improving as retraining becomes less frequent, indicating that frequent retraining often introduces unnecessary variability in forecasts without delivering tangible benefits. These results are statistically validated via statistical tests and indicate that continuous retraining is often not the optimal strategy.
While point forecast accuracy is largely unaffected by reduced retraining, remaining stable or even improving in many cases, probabilistic accuracy shows a slightly higher sensitivity to retraining frequency. Less frequent updates may lead to modest degradations in distributional accuracy, although these effects are generally negligible for low periodic retraining scenarios and remain limited even under no-retraining regimes. Contrary to the common view of a strict trade-off between forecasting accuracy and stability, our results indicate that, in the context of global forecasting models, this tension is largely attenuated. Moreover, in many cases, reducing the retraining frequency does not require sacrificing predictive performance and may even contribute to both more stable and more accurate forecasts. This behavior contrasts with what is often observed in local modeling settings, where model re-estimation can lead to pronounced fluctuations in both accuracy and forecast revisions. For global models, the shared learning across series appears to act as a regularizing mechanism, making them less sensitive to frequent updates and allowing lower retraining frequencies to jointly improve stability while preserving, or even enhancing, also accuracy.
@@
From a practical standpoint, these findings have several implications. First, businesses that rely on large-scale forecasting systems, such as those in retail, energy, and supply chain planning can adopt lower retraining strategies not only to reduce computational resources and the associated costs of forecasting systems, but also to increase operational stability, without compromising on forecast accuracy. Indeed, our results strongly support low, periodic retraining strategies (i.e. every 30 days or every 8 weeks) in global forecasting systems. Finally, these results promote a shift in emphasis from pure accuracy to stability-aware model selection and evaluation, especially in those contexts that are more sensitive to forecast reviews.
@@
Nevertheless, some limitations of the present study should be acknowledged. While the analysis covers a broad range of global models and three state-of-the-art forecasting datasets, exploring the effects of retraining newer methods and different datasets may be a concrete area for future research. Moreover, this work focuses on the stability of global models. A deeper analysis of local approaches could therefore represent a valuable direction for future research. Additionally, the study focuses solely on full retraining schemes and does not explore incremental or online learning strategies, which may offer a different trade-off between accuracy and stability. Future research could extend this work by incorporating such adaptive methods and introducing business-oriented stability metrics that link forecast stability directly to operational cost or risk. Future research should also investigate the interplay between retraining frequency and hyperparameter optimization. Different models may exhibit heterogeneous responses to retraining strategies, depending on their hyperparameter configurations, and a deeper understanding of these interactions could enable the design of even more efficient and robust forecasting strategies. Furthermore, this study employed a single framework for generating probabilistic forecasts, namely Conformal Inference. Although conformalized residuals provide a valid means for quantifying uncertainty, future research could address this limitation by studying probabilistic stability in native probabilistic models, thereby enhancing the robustness and generalizability of the findings. Finally, further theoretical development of stability measures, particularly for probabilistic forecasts, may also contribute to a deeper understanding of how models behave across time and under varying data regimes.
In summary, this work provides strong empirical evidence that frequent retraining is neither necessary nor desirable in many forecasting applications. By explicitly accounting for stability alongside accuracy, forecasting systems can be designed to be more robust and computationally sustainable.
@@
Model Learning rate Iterations Feature fraction
Random Forest 0.03 1200 0.85
XGBoost 0.03 1200 0.85
LightGBM 0.03 1200 0.85
CatBoost 0.03 1200
Table 2: Main hyperparameter settings for machine learning forecasting models.
Model Learning rate Epochs Scaler
MLP 0.001 1000 Robust
LSTM 0.001 1000 Robust
TCN 0.001 1000 Robust
NBEATS 0.001 1000 Robust
NHITS 0.001 1000 Robust
Table 3: Main hyperparameter settings for deep learning forecasting models.
@@
ETS 1.606 1.760 1.925 2.035 2.510 3.041 3.341 3.704 4.113 5.449
ARIMA 1.621 1.776 1.944 2.056 2.545 3.094 3.356 3.804 4.276 5.674
LR 1.578 1.578 1.578 1.578 1.580 1.578 1.581 1.579 1.581 1.584
RF 1.715 1.713 1.713 1.710 1.707 1.699 1.700 1.701 1.701 1.706
XGBoost 18.125 18.115 18.140 18.142 18.144 18.116 18.130 18.026 18.088 18.112
LGBM 15.467 15.452 15.450 15.435 15.413 15.393 15.386 15.282 15.261 15.034
MLP 1.705 1.611 1.676 1.562 1.533 1.546 1.526 1.536 1.529 1.526
TCN 2.280 2.043 1.997 1.815 1.727 1.733 1.699 1.758 1.761 1.730
NBEATSx 1.625 1.521 1.494 1.475 1.468 1.471 1.469 1.473 1.475 1.498
NHITS 1.629 1.548 1.522 1.506 1.493 1.494 1.496 1.496 1.499 1.503
Table 4: M4 RMSSE values for each method and retrain scenario combination.
Method 7 14 21 30 60 90 120 150 180 364
ETS 0.014 0.010 0.008 0.007 0.005 0.004 0.003 0.003 0.002 0.000
ARIMA 0.014 0.010 0.008 0.007 0.005 0.004 0.003 0.003 0.002 0.000
LR 0.013 0.013 0.013 0.013 0.013 0.013 0.013 0.013 0.013 0.013
RF 0.013 0.013 0.013 0.013 0.013 0.013 0.013 0.013 0.013 0.013
XGBoost 0.052 0.052 0.052 0.052 0.052 0.052 0.052 0.052 0.052 0.052
LGBM 0.123 0.123 0.123 0.123 0.123 0.122 0.122 0.122 0.122 0.121
MLP 0.017 0.016 0.017 0.016 0.016 0.016 0.016 0.016 0.016 0.016
TCN 0.021 0.017 0.016 0.015 0.015 0.015 0.015 0.015 0.014 0.015
NBEATSx 0.015 0.015 0.015 0.015 0.015 0.015 0.015 0.015 0.015 0.016
NHITS 0.016 0.015 0.015 0.015 0.015 0.015 0.015 0.015 0.015 0.015
Table 5: M4 sMAPC values for each method and retrain scenario combination.
Method 7 14 21 30 60 90 120 150 180 364
ETS 0.419 0.467 0.513 0.549 0.684 0.864 0.948 1.078 1.230 1.563
ARIMA 0.428 0.475 0.524 0.560 0.701 0.881 0.955 1.111 1.276 1.655
LR 0.394 0.394 0.393 0.395 0.397 0.398 0.397 0.397 0.398 0.400
RF 0.437 0.438 0.437 0.439 0.440 0.440 0.439 0.439 0.440 0.443
XGBoost 2.279 2.288 2.293 2.309 2.337 2.365 2.394 2.405 2.447 2.636
LGBM 2.455 2.473 2.480 2.511 2.564 2.624 2.680 2.707 2.793 3.096
MLP 0.501 0.465 0.470 0.443 0.437 0.439 0.433 0.436 0.437 0.429
TCN 0.652 0.583 0.577 0.531 0.508 0.532 0.506 0.528 0.521 0.507
NBEATSx 0.466 0.426 0.415 0.410 0.409 0.409 0.410 0.407 0.409 0.421
NHITS 0.466 0.435 0.423 0.419 0.414 0.414 0.414 0.412 0.414 0.415
Table 6: M4 SMQL values for each method and retrain scenario combination.
Method 7 14 21 30 60 90 120 150 180 364
ETS 0.278 0.300 0.326 0.351 0.396 0.435 0.449 0.469 0.493 0.603
ARIMA 0.279 0.301 0.328 0.347 0.397 0.439 0.451 0.482 0.496 0.593
LR 0.262 0.262 0.262 0.261 0.263 0.264 0.266 0.266 0.268 0.276
RF 0.351 0.351 0.351 0.352 0.352 0.353 0.354 0.353 0.353 0.357
XGBoost 2.156 2.150 2.154 2.153 2.149 2.142 2.134 2.126 2.130 2.119
LGBM 1.935 1.927 1.928 1.915 1.906 1.904 1.890 1.864 1.861 1.795
MLP 0.371 0.359 0.369 0.346 0.349 0.353 0.354 0.357 0.350 0.360
TCN 0.506 0.397 0.375 0.352 0.323 0.345 0.343 0.332 0.320 0.331
NBEATSx 0.338 0.330 0.330 0.330 0.331 0.333 0.341 0.337 0.341 0.354
NHITS 0.344 0.332 0.329 0.328 0.326 0.326 0.330 0.328 0.331 0.330
Table 7: M4 SMQC values for each method and retrain scenario combination.
Method 7 14 21 30 60 90 120 150 180 364
ETS 0.747 0.751 0.756 0.759 0.769 0.776 0.778 0.774 0.784 0.782
LR 0.759 0.759 0.759 0.759 0.758 0.758 0.758 0.758 0.759 0.758
XGBoost 0.737 0.737 0.737 0.737 0.737 0.737 0.737 0.737 0.736 0.735
LGBM 0.753 0.753 0.753 0.753 0.753 0.753 0.753 0.753 0.753 0.753
CatBoost 0.917 0.929 0.924 0.928 0.921 0.936 0.925 0.930 0.934 0.926
MLP 0.803 0.800 0.800 0.801 0.801 0.802 0.799 0.801 0.801 0.800
TCN 0.844 0.844 0.844 0.844 0.844 0.844 0.844 0.843 0.843 0.841
NBEATSx 0.797 0.796 0.796 0.807 0.797 0.797 0.797 0.797 0.797 0.796
NHITS 0.811 0.808 0.817 0.808 0.808 0.806 0.807 0.807 0.807 0.805
Table 8: M5 RMSSE values for each method and retrain scenario combination.
Method 7 14 21 30 60 90 120 150 180 364
ETS 0.104 0.070 0.054 0.045 0.026 0.018 0.012 0.013 0.007 0.000
LR 0.188 0.188 0.188 0.188 0.189 0.190 0.190 0.191 0.192 0.196
XGBoost 0.075 0.074 0.073 0.073 0.072 0.073 0.072 0.072 0.073 0.073
@@
CatBoost 0.120 0.045 0.060 0.039 0.031 0.022 0.021 0.020 0.020 0.023
MLP 0.499 0.511 0.505 0.519 0.512 0.524 0.510 0.525 0.475 0.552
TCN 0.497 0.208 0.174 0.086 0.055 0.043 0.031 0.017 0.017 0.003
NBEATSx 0.547 0.540 0.527 0.560 0.528 0.535 0.522 0.536 0.546 0.424
NHITS 0.566 0.547 0.557 0.533 0.566 0.536 0.532 0.548 0.511 0.565
@@
ETS 0.249 0.251 0.254 0.256 0.260 0.263 0.264 0.264 0.267 0.288
LR 0.267 0.268 0.269 0.268 0.270 0.271 0.272 0.271 0.272 0.273
XGBoost 0.245 0.246 0.247 0.247 0.249 0.250 0.252 0.251 0.251 0.253
LGBM 0.246 0.247 0.247 0.247 0.248 0.249 0.250 0.249 0.250 0.251
CatBoost 0.277 0.268 0.278 0.272 0.273 0.269 0.265 0.271 0.268 0.267
MLP 0.263 0.264 0.265 0.266 0.269 0.271 0.271 0.272 0.275 0.276
TCN 0.270 0.271 0.272 0.273 0.276 0.278 0.279 0.279 0.281 0.282
NBEATSx 0.262 0.263 0.264 0.268 0.268 0.270 0.271 0.271 0.274 0.275
NHITS 0.266 0.266 0.273 0.268 0.271 0.273 0.274 0.273 0.276 0.278
Table 10: M5 SMQL values for each method and retrain scenario combination.
@@
ETS 1.886 1.910 2.033 1.967 2.032 2.050 2.146 2.398 2.113 2.036
ARIMA 1.935 1.944 1.969 1.996 1.897 1.789 2.010 1.898 1.821 1.830
LR 6.886 6.873 6.925 6.886 6.892 6.888 6.870 6.947 7.037 7.077
RF 1.871 1.873 1.872 1.868 1.867 1.866 1.861 1.863 1.852 1.840
XGBoost 1.898 1.871 1.872 1.865 1.842 1.851 1.809 1.833 1.851 1.797
LGBM 3.540 3.520 3.471 3.282 3.414 3.732 3.415 3.337 3.357 3.543
CatBoost 5.631 5.570 5.694 5.025 4.846 5.359 4.418 4.876 5.212 5.343
MLP 1.554 1.483 1.477 1.472 1.472 1.475 1.480 1.480 1.467 1.455
LSTM 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960
TCN 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960 1.960
NBEATSx 1.704 1.696 1.661 1.561 1.519 1.525 1.544 1.462 1.455 1.456
NHITS 1.705 1.680 1.629 1.556 1.547 1.460 1.528 1.453 1.455 1.451
Table 12: VN1 RMSSE values for each method and retrain scenario combination.
Method 1 2 3 4 6 8 10 13 26 52
ETS 0.156 0.114 0.093 0.070 0.054 0.039 0.032 0.034 0.014 0.000
ARIMA 0.150 0.098 0.080 0.060 0.046 0.034 0.027 0.030 0.012 0.000
LR 0.086 0.084 0.085 0.083 0.083 0.086 0.086 0.085 0.084 0.083
RF 0.093 0.090 0.090 0.090 0.088 0.089 0.089 0.090 0.089 0.090
XGBoost 0.146 0.129 0.125 0.120 0.118 0.115 0.114 0.119 0.116 0.113
@@
CatBoost 0.312 0.197 0.185 0.152 0.134 0.139 0.115 0.117 0.101 0.095
MLP 0.457 0.468 0.467 0.475 0.442 0.483 0.437 0.492 0.468 0.451
LSTM 0.557 0.301 0.202 0.126 0.101 0.044 0.055 0.048 0.032 0.013
TCN 0.599 0.259 0.226 0.177 0.120 0.091 0.052 0.056 0.020 0.013
NBEATSx 0.594 0.592 0.594 0.576 0.576 0.581 0.572 0.580 0.571 0.549
NHITS 0.605 0.587 0.581 0.571 0.531 0.540 0.550 0.527 0.574 0.581
@@
ETS 1.003 1.032 1.077 1.084 1.115 1.133 1.159 1.249 1.196 1.259
ARIMA 0.914 0.936 0.953 0.955 0.974 0.951 1.016 1.021 1.044 1.141
LR 4.270 4.292 4.412 4.378 4.451 4.294 4.295 4.454 4.358 4.669
RF 1.220 1.230 1.232 1.224 1.237 1.254 1.257 1.284 1.313 1.326
XGBoost 1.503 1.496 1.493 1.505 1.483 1.488 1.466 1.451 1.515 1.479
LGBM 2.318 2.292 2.230 2.157 2.230 2.317 2.262 2.362 2.220 2.268
CatBoost 4.203 4.122 4.292 3.970 3.735 4.028 3.549 3.798 3.863 4.192
MLP 0.806 0.800 0.796 0.795 0.803 0.805 0.807 0.819 0.824 0.847
LSTM 0.892 0.895 0.899 0.903 0.910 0.917 0.924 0.935 0.957 1.003
TCN 0.891 0.895 0.899 0.903 0.910 0.917 0.924 0.935 0.957 1.003
NBEATSx 0.991 0.995 0.966 0.897 0.861 0.866 0.893 0.825 0.839 0.875
NHITS 0.992 0.996 0.944 0.894 0.876 0.814 0.880 0.820 0.833 0.862
Table 14: VN1 SMQL values for each method and retrain scenario combination.
Method 1 2 3 4 6 8 10 13 26 52
ETS 0.533 0.460 0.428 0.375 0.347 0.323 0.298 0.296 0.228 0.151
ARIMA 0.463 0.396 0.345 0.304 0.291 0.260 0.255 0.235 0.182 0.132
LR 1.559 1.551 1.545 1.536 1.549 1.525 1.544 1.632 1.674 1.894
@@
Figure 5: RMSSE and sMAPC results for each local method and retrain scenario combination in relative terms with respect to the baseline scenario, r = 7 for the M4 and M5 datasets, and r = 1 for the VN1 dataset.
Figure 6: SMQL and SMQC results for each local method and retrain scenario combination in relative terms with respect to the baseline scenario, r = 7 for the M4 and M5 datasets, and r = 1 for the VN1 dataset.
Figure 7: M4 Friedman-Nemenyi test results based on RMSSE.
Figure 8: M4 Friedman-Nemenyi test results based on sMAPC.
Figure 9: M4 Friedman-Nemenyi test results based on SMQL.
Figure 10: M4 Friedman-Nemenyi test results based on SMQC.
Figure 11: M5 Friedman-Nemenyi test results based on RMSSE.
Figure 12: M5 Friedman-Nemenyi test results based on sMAPC.
Figure 13: M5 Friedman-Nemenyi test results based on SMQL.
@@
Figure 15: VN1 Friedman-Nemenyi test results based on RMSSE.
Figure 16: VN1 Friedman-Nemenyi test results based on sMAPC.
Figure 17: VN1 Friedman-Nemenyi test results based on SMQL.
@@
Figure 19: M4 Friedman-Nemenyi test results for local models.
Figure 20: M5 Friedman-Nemenyi test results for local models.
Figure 21: VN1 Friedman-Nemenyi test results for local models.